Part C Suppose the market for flowers is perfectly competitive. The market dema decimals to 2 decimal places. tsupply curve is qs p-200. In the questions below, round any a. Calculate the perfectly competitive equilibrium q and p
Suppose the government institutes a per-unit tax on consumers equal to 75% of the market price of flowers. Calculate the new equilibrium q and p, consumer incidence, producer incidence, points (original demand and supply curves, new demand and/or s intercepts, slopes, equilibrium q and p, etc.). b. and tax revenue. Graph this market with the tax, labeling all key
c. Now The suppose the market for flowers is a monopoly and there is no s cost curve is cgs) 200+a.. Therefore, the monopolists marginal cost curve is MC(%) 1 and av curve is MR 500- profits. Graph this market, labeling all key points intercepts, slopes, equilibrium q and p. etc.). The marginal revenue erage cost curve is AC(%)-2, 1 Calculate the monopoly equilbrium q and p, and monopoly (demand and marginal cost curves,
Answer #1

2200 5P 3.40 fps すニ293 → R ice Received by seller S500 48 ias tax. 140 し 295 O 6.5 七ー_ , Revenue- Congum, er lncidence: phoduce s beahing. tax ao 6唄5x45 o 15168.75$ Ang 481,25-440 X100: &,0% Now Clo fax 440-a45-X100% 80% = 206, aso) Monopoly 9 ·ccts) MR= 500-V12 5 DO 1 040012 500 2. pr= 250.5 ante MC 8000-998 MR 0-998 1o0 0 -998x 950.5 800-998